Find and correct the errors in the following mathematical statements.
 x+2x+3x = 5x

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

LHS = x+2x+3x= 6x ≠ RHS
The correct statement is x+2x+3x = 6x

Thanks